b.) Calculate the Q-value of a harmonic oscillator carrying a mass of 30 g and is rotating on a circular path of 1.2 m. (2 marks)
1
Expert's answer
2019-10-17T11:12:43-0400

The Q-value of such a harmonic oscillator with the damping coefficient "b" and frequency "n" is


"Q=\frac{\omega_0m}{b}=\frac{2\pi nm}{b}=\frac{0.188n}{b}."